gcl6:

The acetylene tank contains 35.0 mol C2H2, and the oxygen tank contains 84.0 mol O2. How many moles of CO2 are produced when 35.0 mol C2H2 react completely? 67.2 ⇒ 70.0 mol CO2 How many moles of CO2 are produced when 84.0 mol O2 completely react? mol CO2

Narad:

\[2C _{2}H _{2} + 5O _{2} \rightarrow 4CO _{2} + 2H _{2} O\] 2 moles of acetylene will give 4 moles of carbon dioxide. 35 moles of acetylene will give = 4/2*35 =70 moles of carbon dioxide 5 moles of oxygene will give 4 moles of carbon dioxide 84 moles of oxygene will give = 4/5*84= 67.2 moles of carbon dioxide